# VersaTune: Harnessing Vertical Domain Insights for Multi-Ability LLM Supervised Fine-Tuning

### v1

@@ Abstract
Large Language Models (LLMs) exhibit remarkable capabilities in handling multiple tasks across domains due to their emergent properties. These capabilities are further augmented during the Supervised Fine-Tuning (SFT) phase. Despite their potential, existing work mainly focuses on domain-specific enhancements during fine-tuning, the challenge of which lies in catastrophic forgetting of knowledge across other domains. In this study, we introduce VersaTune, a novel data composition framework designed for enhancing LLMs’ overall multi-ability performances during fine-tuning. We categorize knowledge into distinct domains including law, medicine, finance, science, code. We begin with detecting the distribution of domain-specific knowledge within the base model, followed by the composition of training data that aligns with the model’s existing knowledge distribution. During the fine-tuning process, weights of different domains are dynamically adjusted based on their learnable potential and forgetting degree. Experimental results demonstrate that VersaTune achieves significant improvements in multi-domain performance, with a enhancement in comprehensive multi-domain tasks. Additionally, in scenarios where specific domain optimization is required, VersaTune reduces the degradation of performance in other domains by , without compromising the target domain’s training efficacy.
1 Introduction
@@
Large Language Models (LLMs) have become a cornerstone in Artificial Intelligence (AI) (Achiam et al., 2023; Dwivedi et al., 2021; Lewkowycz et al., 2022), particularly for Natural Language Processing tasks (Brown et al., 2020; Devlin, 2018; Radford et al., 2019), reshaping AI research and applications in domains such as law (Cui et al., 2023), medical (Singhal et al., 2023; Thirunavukarasu et al., 2023), finance (Li et al., 2023; Wu et al., 2023), science (Beltagy et al., 2019; Taylor et al., 2022) and code (Liu et al., 2024; Roziere et al., 2023). In order to further incentivize capabilities in these areas, LLMs typically undergo supervised fine-tuning (SFT) stages on specific datasets.
@@
As demonstrated by the robust performances of proprietary LLMs such as GPT-4 (Achiam et al., 2023) and Gemini (Team et al., 2023), LLMs have the potential to master multiple tasks across all specific domains within a single model. However, most existing research on SFT tends to merely concentrate on a single ability of LLMs (Dong et al., 2023; Xu et al., 2024). We try to enhance the overall multitasking performance of LLMs across various domains by optimizing data mixing ratios during the SFT phase:
@@
How to design a data composition strategy during LLMs’ SFT stages that could achieve overall multitasking capabilities?
@@
C1: Catastrophic Forgetting. Given the fundamental differences between tasks across various domains, for multi-domain SFT, the sequential training strategy of specific datasets in multiple phases can easily lead to significant performance drop of prior knowledge, which is well-known as Catastrophic Forgetting (Kaushik et al., 2021; McCloskey & Cohen, 1989). This problem hinders the versatile fine-tuning performance of a model across multiple abilities (De Lange et al., 2021; Dong et al., 2023; Yuan et al., 2022). Therefore, mixing data from different domains is crucial for mitigating catastrophic forgetting during the SFT stage, enhancing the overall performance and adaptability.
@@
C2: Low Efficiency. Existing research on data composition during the SFT phase for LLMs is still in its initial stages, with most strategies based on heuristic rules (Albalak et al., 2024; Dubey et al., 2024; Wang et al., 2023). Such approaches struggle to optimally balance different domains, failing to maximize the overall training effectiveness for multiple abilities. There lacks a well-defined methodology that efficiently enhances the versatile capabilities of LLMs across multiple domains during the SFT stage.
@@
C3: Low Flexibility in Domain Expansion. Existing SFT approaches for specific domain abilities typically pre-determine the proportions of different datasets according to prior experience (Azerbayev et al., 2023; Roziere et al., 2023). Such strategies lack the flexibility to dynamically adjust the data mixing ratios of domains during the current phase based on real-time feedback from the LLMs, which hampers the ability to minimize the loss of performances in other domains as the LLMs undergo specialized training.
@@
To address these challenges, we introduce VersaTune, a novel data composition framework designed for enhancing models’ overall performances across different vertical domains during fine-tuning. We first detect the proportion distribution of domain knowledge within the target model (Section 2.2), followed by data composition based on the existing distribution for multi-ability enhancement (Section 3.2) as well as flexible domain expansion (Section 3.3). VersaTune exhibits the following properties:
@@
Efficient. VersaTune employs distribution consistency training of the domain knowledge proportion during models’ fine-tuning stage, providing an efficient strategy for enhancing versatile capabilities (C2).
@@
Flexible. VersaTune can be flexibly adapted to scenarios that expand performance on specific domain tasks while minimizing the degradation of the model’s capabilities in non-target domains (C1, C3).
@@
Robust. The data composition strategy demonstrates significant improvements after fine-tuning in open-source models with parameter sizes ranging from 7B-14B, adding to the effectiveness of VersaTune .
@@
We evaluate our VersaTune on 12 downstream tasks across vertical domains, including law, medicine, finance, science, and code, as well as comprehensive datasets, each of which features two benchmarks. The results show that VersaTune can achieve an improvement of in versatile fine-tuning across multiple domains. Furthermore, when focusing on specific-domain fine-tuning, VersaTune maintains the training effectiveness in the target domain while reducing the performance degradation in other non-target domains by .
2 VersaTune: Detecting Knowledge Distribution
@@
In this section, we present a domain mixing strategy for fine-tuning a LLM that possesses a comprehensive multi-task capability (Section 2.1). This approach is designed to align with the inherent domain knowledge distribution of the base model. Subsequently, we describe the method for detecting domain knowledge proportion of the base model, which is crucial for informing the fine-tuning process (Section 2.2).
2.1 Knowledge Consistency Training
@@
Statement 1 An LLM fine-tuned with domain-specific data proportions that align with its pre-trained output distributions will exhibit enhanced and balanced performance across these domains, compared to a model fine-tuned with a non-matching data distribution. Formally, the relationship can be represented as:
@@
The statement is based on the fact that the model has developed an understanding of language features and knowledge specific to various domains during pretraining. Maintaining the same distribution of knowledge during fine-tuning allows the model to build on this pre-existing knowledge, enhancing learning efficiency and robustness.
2.2 Domain Knowledge Distribution Detection
Drawing on prior research into knowledge identification methods (Gekhman et al., 2024; Zhao et al., 2023b) and training data inference strategies for LLMs (Ding et al., 2022; Hayase et al., 2024), we propose a structured approach to efficiently detect domain knowledge based on statistics.
Assuming the data corpus encompasses distinct domains, as shown in Algorithm 1, we first prompt the base model poised for fine-tuning with the Beginning of Sequence (<BOS>) tokens to generate a set of data entries (Line 3). Subsequently, we employ a proprietary LLM to infer the probabilities that these entries belong to each domain (Line 5-7). We then calculate a weighted average of the probability distributions for all data across these domains, thereby deriving the domain knowledge distribution of the current base model (Line 9). To ensure statistical robustness, the process is iteratively conducted times, and we use the mean of these iterations as the estimated result for the domain knowledge distribution.
@@ 2.2 Domain Knowledge Distribution Detection
3 VersaTune: Fine-Tuning Multi-Ability LLMs Efficiently
Having detected the distribution of domain knowledge within the base model, we will now utilize these findings to guide our multi-ability SFT process. The approaches are designed to enhance the overall performance of the fine-tuned model across a spectrum of multi-domain tasks (Section 3.2), as well as to facilitate the flexible expansion of capabilities in specific domains (Section 3.3).
Setting. Our goal is to construct a composite dataset covering specific domains, which can be denoted as with each tuple representing a domain and its corresponding proportion, such that training a model on could achieve overall lower loss on a uniformly distributed composite target validation dataset or meet the flexible domain expansion while preserving the performance of other domains. The specialized capabilities of LLMs are ultimately measured using downstream tasks related to different domains (e.g., FinBen (Xie et al., 2024a) for financial performances).
3.1 Preliminary: Learnable Potential and Forgetting Degree of Domain Knowledge

### v3

# VersaTune: An Efficient Data Composition Framework for Training Multi-Capability LLMs

@@ Abstract.
Large-scale pretrained models, particularly Large Language Models (LLMs), have exhibited remarkable capabilities in handling multiple tasks across domains due to their emergent properties. These capabilities are further augmented during the Supervised Fine-Tuning (SFT) phase. Despite their potential, existing work mainly focuses on domain-specific enhancements during fine-tuning, the challenge of which lies in catastrophic forgetting of knowledge across other domains. In this study, we introduce VersaTune, a novel data composition framework designed for enhancing LLMs’ overall multi-ability performances during training. We categorize knowledge into distinct domains including law, medicine, finance, science, code, etc. We begin with detecting the distribution of domain-specific knowledge within the base model, followed by the training data composition that aligns with the model’s existing knowledge distribution. During the training process, domain weights are dynamically adjusted based on their learnable potential and forgetting degree. Experimental results demonstrate that VersaTune achieves significant improvements in multi-domain performance, with an enhancement in comprehensive multi-domain tasks. Additionally, in scenarios where specific domain optimization is required, VersaTune reduces the degradation of performance in other domains by , while preserving the target domain’s training efficacy.
PVLDB Reference Format:
Keer Lu, Keshi Zhao, Zheng Liang, Da Pan, Shusen Zhang, Xin Wu, Zenan Zhou, Guosheng Dong, Bin Cui, Wentao Zhang. PVLDB, 18(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX
† Corresponding Author.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and other artifacts will be released after the paper is accepted.
1. Introduction
@@
In recent years, large-scale pretrained models with enormous pretraining tokens and parameters, such as Transformer models, have become a cornerstone in Artificial Intelligence (AI) (Achiam et al., 2023; Dwivedi et al., 2021; Lewkowycz et al., 2022), particularly for Natural Language Processing tasks (Brown et al., 2020; Devlin, 2018; Radford et al., 2019), reshaping AI research and applications in domains such as law (Cui et al., 2023), medicine (Singhal et al., 2023; Thirunavukarasu et al., 2023), finance (Li et al., 2023b; Wu et al., 2023), science (Beltagy et al., 2019; Taylor et al., 2022) and code (Liu et al., 2024; Roziere et al., 2023). The success of Transformer models, such Large Language Models (LLMs), stems from their capabilities to automatically learn and distill hierarchical data representations, making them highly effective for complex tasks (Nie et al., 2023). In order to further enhance such abilities across these areas, LLMs typically undergo the supervised fine-tuning (SFT) stages on domain-specific datasets.
Data plays a fundamental role throughout the training phases of LLMs. While the significance of model architectures is widely acknowledged, data-centric approaches hold substantial importance (Wang et al., 2023), which emphasizes the quality, organization, and management of data as the crucial factor for improving model performance (Miao et al., 2024; Fernandez et al., 2023; Whang et al., 2023). Efficient data management, especially in constructing an appropriately tailored training dataset with suitable data composition, is critical for synergizing the model’s performance and ability to generalize across different tasks and domains during both pretraining and supervised fine-tuning stages (Wang et al., 2023; Kumar et al., 2017).
@@
As demonstrated by the robust performances of state-of-the-art LLMs such as GPT-4 (Achiam et al., 2023), PaLM (Anil et al., 2023) and Gemini (Team et al., 2023), LLMs have the potential to master multiple tasks across all specific domains within a single model. However, most existing research on supervised fine-tuning tends to merely concentrate on a single ability of LLMs (Dong et al., 2023; Xu et al., 2024), with the multi-task performance on composite data of essentially different downstream tasks being less studied. We try to enhance the overall multitasking performance of LLMs across various domains by optimizing data mixing ratios during training:
@@
How to design a data composition strategy that could achieve overall multitasking capabilities?
@@
C1: Catastrophic Forgetting. Given the fundamental differences between tasks across various domains, for multi-domain SFT, the sequential training strategy of specific datasets in multiple phases can easily lead to significant performance drop of prior knowledge, which is well-known as Catastrophic Forgetting (Kaushik et al., 2021; McCloskey and Cohen, 1989), as illustrated in Table 1 and Figure 2. This problem hinders the versatile fine-tuning performance of a model across multiple domains (De Lange et al., 2021; Dong et al., 2023; Yuan et al., 2022). Therefore, mixing data from different domains is crucial for mitigating catastrophic forgetting during the SFT stage, enhancing the overall performance and adaptability.
@@
C2: Low Efficiency. Existing research on data composition during the supervised fine-tuning phase for LLMs is still in its initial stages, with most strategies based on heuristic or manually determined rules (Albalak et al., 2024; Dubey et al., 2024; Wang et al., 2023). One of the common baselines is defining domain weights referring to the natural domain sizes, which weights all the individual data points equally. Such approaches struggle to optimally balance different domains, failing to maximize the overall training effectiveness for multiple abilities. There lacks a well-defined methodology that efficiently enhances the versatile capabilities of LLMs across multiple domains during the SFT stage.
@@
C3: Low Flexibility in Domain Expansion. Existing SFT approaches for specific domain abilities typically pre-determine the proportions of different datasets according to prior experience (Azerbayev et al., 2023; Roziere et al., 2023). Such strategies lack the flexibility to dynamically adjust the data mixing ratios of different domains during the training process, which does not allow for real-time feedback from the LLMs in the current phase to inform and optimize the data composition. This static approach hinders the minimization of performance loss in other domains as LLMs undergo specialized training.
@@
To address these challenges, we introduce VersaTune, a novel data composition framework designed for enhancing models’ overall performances across different vertical domains during supervised fine-tuning. We first detect the proportion distribution of domain knowledge within the target model (Section 3.1), followed by data composition based on the existing distribution for multi-ability enhancement (Section 3.2.2) as well as flexible domain expansion (Section 3.2.3). VersaTune exhibits the following properties:
@@
Efficient. VersaTune employs distribution consistency training of the domain knowledge proportion during models’ fine-tuning stage, providing an efficient data composition strategy for enhancing versatile capabilities (for C2).
@@
Flexible. VersaTune can be flexibly adapted to scenarios that expand performance on specific domain tasks while minimizing the degradation of the model’s capabilities in non-target domains (for C1, C3).
@@
Robust. The data composition strategy demonstrates significant improvements after fine-tuning in open-source models with parameter sizes ranging from 7B-14B, adding to the effectiveness of VersaTune (for C1, C2 and C3).
In summary, our contributions are as follows:
Challenges in Multi-Capability Training. Through a quantitative analysis of the performance losses in non-target domains resulting from existing training methods, we emphasize the challenge of data composition in multi-domain task training (Section 2).
Multi-Capability Data Composition Framework. We introduce VersaTune, a data composition framework that leverages the intrinsic domain knowledge distribution of the model to proportion training data. VersaTune is designed to enhance overall performance across multiple domains while also providing flexibility for targeted improvements in specific domains (Section 3).
@@
Performance and Effectiveness. Our comprehensive evaluations across domains demonstrate that VersaTune can achieve an improvement of in versatile fine-tuning for multiple domains. Furthermore, when focusing on specific-domain fine-tuning, VersaTune maintains the training effectiveness in the target domain while reducing the performance degradation in other non-target domains by (Section 5).
2. Preliminary
2.1. Transformer Models
Transformers were initially introduced to address sequence modeling and transduction challenges, including language modeling and machine translation (Vaswani, 2017). The core of each Transformer layer consists of self-attention mechanisms and point-wise feed-forward networks. Self-attention network enables the model to assess the importance of various words in a sequence regardless of their positions (Zhang et al., 2019). It calculates a score for each word, indicating the degree of attention to be directed towards other sequence parts during encoding. This process involves query, key, and value weight sets that are optimized through training. Feed-forward network (FFN) applies the same linear transformation with a ReLU activation to each position individually. The FFN’s role is to independently transform each position’s representation, introducing non-linearity into the model and allowing it to learn complex functions (Miao et al., 2022). This component complements the self-attention mechanism by processing each input position in isolation, thus enhancing the model’s adaptability to diverse data types and tasks. These operations, predominantly dense algebraic computations such as matrix multiplications, lead to significant computational expenses and memory demands.
| Domain of Training Dataset | Training Step | Variations in Comprehensive Domains () | Sum. () | |||||
| (Epoch) | Law | Medicine | Finance | Science | Code | Other | ||
| Law | 1 | - | 18.82 | 14.71 | 11.76 | 11.18 | 5.00 | 32.05 |
| 2 | - | 12.65 | 30.59 | 4.41 | 5.29 | 11.76 | 3.52 | |
| 3 | - | 17.94 | 12.35 | 8.82 | 23.53 | 5.00 | 42.94 | |
| 4 | - | 5.00 | 2.06 | 31.18 | 21.76 | 12.65 | 72.65 | |
| Medicine | 1 | 10.29 | - | 3.82 | 24.12 | 19.41 | 7.35 | 16.75 |
| 2 | 18.82 | - | 6.47 | 40.00 | 7.94 | 17.65 | 10.88 | |
| 3 | 22.35 | - | 8.82 | 7.94 | 19.12 | 10.00 | 52.35 | |
| 4 | 27.94 | - | 11.76 | 2.35 | 21.76 | 12.65 | 76.46 | |
| Finance | 1 | 20.59 | 7.94 | - | 10.29 | 12.65 | 6.47 | 16.76 |
| 2 | 18.24 | 9.71 | - | 9.41 | 5.29 | 8.82 | 4.41 | |
| 3 | 23.53 | 9.41 | - | 17.35 | 14.71 | 7.94 | 25.88 | |
| 4 | 5.00 | 9.12 | - | 20.29 | 12.94 | 21.76 | 59.11 | |
| Science | 1 | 10.29 | 17.06 | 3.82 | - | 4.71 | 7.35 | 9.11 |
| 2 | 11.47 | 12.35 | 4.71 | - | 5.88 | 12.94 | 22.65 | |
| 3 | 21.76 | 7.94 | 8.82 | - | 4.41 | 10.00 | 37.05 | |
| 4 | 27.94 | 2.35 | 11.47 | - | 12.65 | 12.59 | 62.30 | |
| Code | 1 | 3.82 | 7.35 | 17.35 | 9.12 | - | 7.29 | 11.99 |
| 2 | 9.71 | 6.47 | 7.94 | 5.29 | - | 6.18 | 25.01 | |
| 3 | 22.35 | 8.82 | 14.12 | 7.94 | - | 10.02 | 47.37 | |
| 4 | 26.18 | 7.06 | 8.82 | 3.24 | - | 22.65 | 67.95 | |
2.2. LLMs Training
Large Language Models (LLMs) are Transformer-based models containing an extensive parameter count in the billions or greater, which are trained on a vast scale of datasets (Zhao et al., 2023b). The training process of LLMs generally involves the pretraining and fine-tuning stages. We have outlined several concepts about LLMs training.
2.2.1. Pretraining
Large Language Models (LLMs) establish basic knowledge abilities, including language understanding and text generation, during the pretraining stage (Brown et al., 2020). In this stage, LLMs engage in unsupervised training through the processing of extensive raw text corpora, thereby enhancing their capabilities in language modeling. For a given sequence , the typical task for LLMs involves the prediction of the subsequent token given the preceding tokens as contextual input. The goal is to maximize the likelihood function presented in Equation 1:
Beginning of Sequence (¡BOS¿) During the above process, the Beginning of Sequence (¡BOS¿) token plays an important role, which serves as a signal to the model that the input sequence is starting (Brown et al., 2020; Li et al., 2023a; Wu et al., 2025). It can be thought of as a special marker that indicates the start of a new sequence, allowing the model to reset its context and begin processing a new piece of text. In the context of pretraining, the ¡BOS¿ token is used to initialize the input to the model, and it can be concatenated with the actual text data to form the input sequence. This token helps the model to differentiate between the start of a new input and the continuation of an existing one. It is particularly crucial in tasks where the model needs to generate text or understand the beginning of a new sentence or document, which helps the model to learn the boundaries of text sequences and to better model the statistical properties of the language data it is trained on. The use of ¡BOS¿ tokens, along with other special tokens like ¡EOS¿ (End of Sequence), helps the model to learn the boundaries of text sequences and to better model the statistical properties of the language data it is trained on.
2.2.2. Supervised Fine-Tuning
The Supervised Fine-Tuning (SFT) stage involves further training on the pretrained model to refine it task-solving capabilities and ensure greater alignment with human instructions across multiple domains (Zhao et al., 2023b; Nie et al., 2023). During the SFT phase, it is a typical practice to employ datasets specific to a particular domain for the fine-tuning of LLMs, which may lead to a decline in performance on non-target domains, a phenomenon commonly referred to as catastrophic forgetting. We conducted experiments on open-sourced models including LLaMA (Dubey et al., 2024; Touvron et al., 2023a, b) and Qwen (Bai et al., 2023; Yang et al., 2024) series to assess how the model’s proficiency in other domains changes when fine-tuned with data from a single domain, as illustrated in Table 1 and Figure 2. We have regulated the number of training instances per epoch to a fixed count of 10,000. More details about training and evaluation settings can be found in Section 5.1. Our findings indicate that when a model is trained exclusively with data from a single domain, its performance on tasks from other domains tends to degrade progressively over the course of training.
While recent research has delved into exploring fine-tuning methods for multi-task enhancement (Dong et al., 2023; Sanh et al., 2021), they are still in their early stages. However, as shown by proprietary models such as GPT-4 (Achiam et al., 2023) and Gemini (Team et al., 2023), which exhibit outstanding multi-task performance, improving a model’s versatile capabilities across various domains during the SFT phase is crucial. Therefore, our work systematically investigates methods to enhance multi-domain performance during the SFT stage to bridge this gap.
3. VersaTune
In this section, we introduce VersaTune, a novel data composition framework designed for multi-capability training. This framework aims to effectively compose data from multiple domains and optimize the data composition during training. Figure 3 presents the workflow of VersaTune, which generally contains two phases.
3.1. Phase 1: Knowledge Distribution Detection
@@
Here, we first present a domain mixing strategy for fine-tuning a LLM that possesses a comprehensive multitask capability (Section 3.1.1). This approach is designed to align with the inherent domain knowledge distribution within the base model waiting for subsequent training. Following this, we describe the method for detecting domain knowledge proportion of the base model, which is crucial for informing the fine-tuning process (Section 3.1).
3.1.1. Knowledge Consistency Training
@@
Statement 1 (Knowledge Consistency Training). An LLM fine-tuned with domain-specific data proportions that align with its pretrained output distributions will exhibit enhanced and balanced performance across these domains, compared to a model fine-tuned with a non-matching data distribution. Formally, the relationship can be represented as:
@@
The rationale behind this statement is rooted in the observation that during the pretraining phase, LLMs develop a general understanding of language features and domain-specific knowledge. By maintaining the same distribution of knowledge during fine-tuning, the model can build upon this pre-existing knowledge, thereby enhancing learning efficiency and robustness.
3.1.2. Domain Knowledge Distribution Detection
Drawing on prior research into knowledge identification methods (Gekhman et al., 2024; Zhao et al., 2023a) and training data inference strategies for LLMs (Ding et al., 2022; Hayase et al., 2024), we propose a structured approach to efficiently detect domain knowledge based on statistics. The process involves the generation of textual outputs from the base model waiting to be fine-tuning, followed by their classification into predefined domains referring to a proprietary model . The above process is repeated multiple times to ensure statistical robustness.
Assuming the data corpus encompasses distinct domains, as shown in Algorithm 1, we first prompt the base model poised for fine-tuning with the Beginning of Sequence (<BOS>) tokens to generate a set of data entries (Line 3). Subsequently, we employ a proprietary model to infer the probabilities that these entries belong to each domain (Line 5-7). We then calculate a weighted average of the probability distributions for all data across these domains, thereby deriving the domain knowledge distribution of the current base model (Line 9). To ensure statistical robustness, the process is iteratively conducted times, and we use the mean of these iterations as the estimated result for the domain knowledge distribution.
@@ 3.1.2. Domain Knowledge Distribution Detection
3.2. Phase 2: Fine-Tuning Multi-Ability LLMs Efficiently
Having detected the distribution of domain knowledge within the base model, we will now utilize these findings to guide our multi-ability SFT process. The approaches are designed to enhance the overall performance of the fine-tuned model across a spectrum of multi-domain tasks (Section 3.2.2), as well as to facilitate the flexible expansion of capabilities in specific domains (Section 3.2.3).
Setting. Our goal is to construct a composite dataset covering specific domains, which can be denoted as with each tuple representing a specific domain and its corresponding proportion, such that training a model on dataset could achieve overall lower loss on a uniformly distributed composite target validation dataset or meet the flexible domain expansion while preserving the performance of other domains. The specialized capabilities of LLMs are ultimately measured using downstream tasks related to different domains (e.g., FinBen (Xie et al., 2024a) for financial performances).
3.2.1. Learnable Potential and Forgetting Degree of Knowledge